the complement of an angle is six less than twice the measure of an angle itself. find the angle and its complement.
Answer #1
Hello! I will just show you how to solve this. Representation: let x=angle, 2x-6= the complement. Equation: x+(2x-6)=90 Solution: x+2x-6=90 3x-6=90 3x=96 x=32 Final
answered by: gggg
Answer #2
: The angle is 32 and the complement is 58.
answered by: Tools(plz help its really urgent)